The number of natural numbers less that or eual to 2985984 which are neither perfect squares nor perfect cubes are:


a) 2984124 b) 2984244 c) 2959595 d) none.


please explain to me how to solve this problem.

Grade:

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

2985984=1728
32985984=144

so the number of perfect squares less than equal to 2985984 = 1728
so the number of perfect cubes less than equal to 2985984 = 144

Number of numbers which are both cubes and perfect squares=12


Numbers which are perfect squares or cubes=1728+144-12=1716+144=1860

hence the numbers which are neither squares nor cubes is
2985984-1860=2984124.

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

sagarsingh24.iitd@gmail.com